\begin{thebibliography}{45}


\bibitem{GasparRahman}  
Gasper, G., Rahman, M.:
{Basic Hypergeometric Series} 
(with a Foreword by Richard Askey), Encyclopedia of 
Mathematics and Its Applications, Volume {35};
Cambridge University Press: Cambridge, UK; New York, NY, USA; 
Port~Chester, NY, USA; Melbourne, Australia; Sydney, Australia, (1990); 
see also 2nd ed., Encyclopedia of Mathematics and 
Its Applications, Volume {96};  Cambridge 
University Press: Cambridge, UK; London, UK; 
New York, NY, USA, (2004) 
 
\bibitem{Slater} 
Slater, L.J.: { Generalized Hypergeometric Functions}; 
Cambridge University Press: Cambridge, UK; London, UK; 
New York, NY, USA, (1966)
 
\bibitem{SrivastavaKarlsson} 
Srivastava, H.M., Karlsson, P.W.:  { Multiple Gaussian 
Hypergeometric Series}; Halsted Press (Ellis Horwood Limited), 
Chichester, UK; John Wiley and Sons: New York, NY, USA; Chichester, UK; 
Brisbane, Australia; Toronto, Canada, (1985)
 
\bibitem{Koekock} 
Koekock, R.,  Swarttouw, R.F.:  { The Askey-Scheme of Hypergeometric 
Orthogonal Polynomials and its $q$-Analogue}; Report No. 98-17;
Delft University of Technology: Delft, The Netherlands, (1998)
 
\bibitem{HMS-IMAJAM1983-1984a} 
Srivastava, H.M.:  {Certain $q$-polynomial expansions for functions of 
several variables. I}. {IMA J. Appl. Math.}, {\bf 30}, 315--323 (1983) 
 
\bibitem{HMS-IMAJAM1983-1984b} 
Srivastava, H.M.:  {Certain $q$-polynomial expansions for functions of 
several variables. II}.
\emph{IMA J. Appl. Math.}  {\bf 33}, 205--209 (1984)
 
\bibitem{HMS-ISTT2020}
Srivastava, H.M.:  {Operators of basic $($or $q$-$)$ calculus and 
fractional $q$-calculus and their applications in geometric function 
theory of complex analysis}. {Iran. J. Sci. Technol. Trans. A$:$ Sci.}, {\bf 44}, 327--344 (2020) 

\bibitem{Srivastava2012} 
Srivastava, H.M., Choi, J.:  {Zeta and $q$-Zeta
Functions and Associated Series and Integrals}; 
Elsevier Science Publishers, Amsterdam, The Netherlandsl 
London, UK; New York, NY, USA, (2012)
 
\bibitem{GEAndrews1986} 
Andrews, G.E.: { $q$-Series$:$ Their Development and 
Applications in Analysis$,$ Number
Theory$,$ Combinatorics$,$ Physics and Computer Algebra}; 
CBMS Regional Conference Lecture Series; 
American Mathematical Society: Providence, Rhode Island, 
  {\bf 66} (1986)
 
\bibitem{HMS-JC-SA2020}
Srivastava, H.M,  Cao, J.,  Arjika, S.:  A Note on Generalized $q$-Difference Equations 
and Their Applications Involving $q$-Hypergeometric~Functions.  {Symmetry} {\bf 12}, Article ID 1816, 1-16 (2020)

\bibitem{14}
 Chen, W.Y.C.,   Liu, Z.-G.: {Parameter augmenting for basic hypergeometric series,I}.  in Mathematical Essays in Honor of Gian-Carlo Roto, B.E Sagan and R.P Stanley, eds., Birkauser, Basel,  111--129 (1998)

\bibitem{15}
Chen, W.Y.C.,   Liu, Z.-G.: {Parameter augmenting for basic hypergeometric series,II}.  J. Combin. Theory, Ser. A \textbf{80}, 175--195  (1997)

\bibitem{SrivastaAbdlhusein} 
Srivastava, H.M., Abdlhusein, M.A.:  {New forms of the Cauchy 
operator and some of their applications}. 
{Russian J. Math. Phys. } {\bf 23}, 124--134 (2016)

\bibitem{11}
 Roman, S.: {The theory of the umbral calculus I}, J. Math. Anal. Appl. \textbf{87},  58--115 (1982)

\bibitem{18}
 Liu, Z.-G.: {Two $q$-difference equations and $q$-operator identities}. ~J. Differ. Equ. Appl. {\bf 16}, 1293-1307  (2010)

\bibitem{1}
Al-Salam, W.A.,  Carlitz,  L.: {Some orthogonal $q$-polynomials}. Math. Nachr. \textbf{30}, 47--61 (1965)


 \bibitem{5}  Koekoek, R.  Lesky, P. A.,   Swarttouw, R. F.:   {
 Hypergeometric Orthogonal Polynomials and Their $q$-Analogues},    {Springer Monographs in Mathematics}, {Springer-Verlag Berlin Heidelberg}, (2010)

\bibitem{cao-niu2016}
Cao, J., Niu, D.-W.: {A note on $q$-difference equations for Cigler's polynomials}.  J. Difference Equ. Appl. \textbf{22},  1880--1892  (2016)

\bibitem{cao-sri-liu}
Cao, J., Srivastava, H.M., Liu,  Z.-G.: {Some iterated fractional $q$-integrals and their applications}.  Fract. Calc. Appl. Anal. \textbf{21}, 672--695 (2018)

\bibitem{cao}
Cao, J.: {A note on $q$-difference equations for Ramanujan's integrals}. Ramanujan J. \textbf{48},  63--73  (2019)


\bibitem{wang2012}
 Wang, M.: {An identity from the Al-Salam--Carlitz polynomials}. Math. AEterna \textbf{2},  185--187  (2012)

\bibitem{wang2013}
 Wang, M.: {A transformation for the Al-Salam--Carlitz polynomials}. Ars Combin. \textbf{112}, 411--418  (2013)


\bibitem{17}
Liu, Z.-G.: {Some operator identities and $q$-series transformation formulas}. Discret. Math. \textbf{265},  119--139  (2003)


\bibitem{fang2013}
Fang, J.-P.: {Remarks on homogeneous Al-Salam and Carlitz polynomials}.  J. Math. 2014, Art. ID 523013, 12 pp (2014)

\bibitem{fang2014}
Fang, J.-P.: {$q$-Difference equation and q-polynomials}.  Appl. Math. Comput. \textbf{248}, 550--561 (2014)

\bibitem{fang2015}
Fang, J.-P.: {Remarks on a generalized $q$-difference equation}.  J. Difference Equ. Appl. \textbf{21},  934--953 (2015)

\bibitem{jia2014}
Jia, Z.: {Two new $q$-exponential operator identities and their applications}. J. Math. Anal. Appl. \textbf{419}, 329--338  (2014)

\bibitem{6}
Srivastava, H.M., Arjika,   S., Sherif Kelil,  A.: {Some homogeneous $q$-difference operators and
the associated generalized Hahn polynomials}. Appl. Set-Valued Anal. Optim. \textbf{1}, 187--201  (2019)

\bibitem{16}
Liu, Z.-G.:  {An extension of the non-terminating ${}_6\psi_5$ summation and the Askey-wilson polynomials}. J. Differ. Equ. Appl. \textbf{17}, 1401--1411   (2011)

\bibitem{10}
Arjika, S.:   {$q$-Difference equation for   homogeneous $q$-difference operators and their applications}. {J. Differ. Equ.~Appl.} {\bf 26}, 987--999 (2020)


\bibitem{19}
Liu, Z.-G.:  {On the $q$-partial differential equations and $q$-series, In: The legacy of Srinivasa Ramanujan}, Ramanujan Mathematical Society Lecture Note Series, Mysore {\bf 20},  213--250, (2013)

\bibitem{7}
Taylor, J.: {Several complex variables with connections to algebraic geometry and lie groups}, Graduate Studies in 
Mathematics, American Mathematical Society, Providence, {\bf 46}, (2002)

\bibitem{9}
 Gunning, R.: {Introduction to Holomorphic Functions of Several Variables. In: Function theory} {\bf 1}, Wadsworth and Brooks/Colc, Bclmont, (1990)

\bibitem{3}
Malgrange, B.: {Lectures on the Theory of Functions of Several Complex Variables}, Springer, Berlin (1984)

\bibitem{13}
Chen, W.Y.C., Fu, A.M.,  Zhang,  B.: {The homogeneous $q$-difference operator}. Adv. App. Math. \textbf{31}, 659--668  (2003)
 
\bibitem{Cao-Srivastava2013}
Cao, J.; Srivastava, H.M.: {Some $q$-generating functions of 
the Carlitz and Srivastava-Agarwal types associated 
with the generalized Hahn polynomials and the 
generalized Rogers-Szeg\"{o} polynomials}. 
{Appl. Math. Comput.}  {\bf 219}, 8398--8406 (2013) 

\bibitem{HMS-C-W2020}
Srivastava, H.M., Chaudhary, M.P., Wakene, F.K.:  {A family of 
theta-function identities based upon $q$-binomial theorem and 
Heine's transformations}. {Montes Taurus J. Pure Appl. Math.} {\bf 8}, 918 (2020) 
%
%
%\bibitem{Gasper-Rahman}
%Gasper, G.,  Rahman, M.: {Basic Hypergeometric Series},~Camberdge:~Camberdge Univ Press, (2004)

\bibitem{20}
Liu, Z.-G.: {$q$-Difference equation and the Cauchy operator identities}. J. Math. Anal. Appl. \textbf{359}, 265--274  (2009)


\bibitem{8}
Milne,  S.C.: {Balanced ${}_3\phi_2$ summation theorenss for $U(n)$ basic hypergeometric series}. Adv. Math. \textbf{131}, 93--187 (1997)


\bibitem{12}
Wang, M.: {$q$-Integral representation of the Al-Salam--Carlitz polynomials}. Appl. Math. Lett. \textbf{22}, 943--945  (2009)


\bibitem{2}Askey, R.: {Two integrals of Ramanujan}.  Proc. Am. Math. Soc. \textbf{85}, 192--194 (1982)
 
\bibitem{HMS-JAEC2019}
Srivastava, H.M.:  The Zeta and related functions: 
Recent developments. {J. Adv. Engrg. Comput.}  {\bf 3}, 329--354 (2019)
 
\bibitem{HMS-PIMM2019}
Srivastava, H.M.:  Some general families of the 
Hurwitz-Lerch Zeta functions and their applications: 
Recent developments and directions for further 
researches. { Proc. Inst. Math. Mech. Nat. 
Acad. Sci. Azerbaijan}  
{\bf 45}, 234--269 (2019)
 
\bibitem{HMS-Symmstry2020}
Shafiq, M., Srivastava, H.M.,  Khan, N., Ahmad, Q.Z., 
Darus, M.,  Kiran, S.: 
An upper bound of the third Hankel determinant for a subclass of 
$q$-starlike functions associated with $k$-Fibonacci numbers.  
{Symmetry} {\bf 12}, Article ID 1043, 1--17 (2020)
 
\bibitem{HMS-Sama2020} 
Srivastava, H.M., Arjika, S.: {Generating functions for some 
families of the generalized Al-Salam-Carlitz $q$-polynomials}. 
{Adv. Differ. Equ.}  {\bf 2020}, 498 (2020).
 
\end{thebibliography}

\documentclass[11pt,reqno]{amsart}
\usepackage{graphicx}
\usepackage{amsfonts}
\usepackage{mathrsfs}
 \usepackage{times}
 \usepackage{mathptmx}
\usepackage{amsmath}
\usepackage{txfonts}
\usepackage{amssymb}
\usepackage{amsmath,amsthm}
\usepackage{amsfonts}
\usepackage[body={7.5in,8.75in}, top=1.2in, right=0.9in,left=0.9in]{geometry}
\usepackage[hypertex,colorlinks=true]{hyperref}
\allowdisplaybreaks[4] \numberwithin{equation}{section}

\newcommand{\abs}[1]{\left\lvert#1\right\rvert}
\newtheorem{thm}{Theorem}

\newtheorem{cor}[thm]{Corollary}
\newtheorem{lem}[thm]{Lemma}
\newtheorem{rem}[thm]{Remark}
\newtheorem{pro}[thm]{Proposition}

\DeclareMathOperator{\td}{d\mspace{-0.5mu}}
\def\squarebox#1{\hbox to #1{\hfill\vbox to #1{\vfill}}}

\begin{document}

\title[On generalized $q$-difference equations  for general Al-Salam--Carlitz polynomials]{A note on generalized $q$-difference equations  for general Al-Salam--Carlitz polynomials}
\author{Jian Cao$^1$, Binbin Xu$^2$ and Sama Arjika${}^{3,*}$ }
\dedicatory{\textsc}
\thanks{${}^{3,*}$Address for correspondence: Department of Mathematics and Informatics,   University of Agadez, Niger. ${}^1$Department
of Mathematics, Hangzhou Normal University, Hangzhou City, Zhejiang Province, 311121, China. ${}^2$Department
of Mathematics, Hangzhou Normal University, Hangzhou City, Zhejiang Province, 311121, China.}
\thanks{Email: 21caojian@hznu.edu.cn,418272337@qq.com, rjksama2008@gmail.com.}

\keywords{$q$-Difference equation; $q$-Difference operator; Al-Salam--Carlitz polynomials; Generating functions; Ramanujan's integral.}

\thanks{2010 \textit{Mathematics Subject Classification}. 05A30,
11B65, 33D15, 33D45, 33D60, 39A13, 39B32.}

\begin{abstract}
In this paper, we  deduce the generalized $q$-difference equations  for general Al-Salam--Carlitz polynomials and generalize  Arjika's recently results [$q$-difference equation for homogeneous $q$-difference operators and their applications, J. Differ. Equ.  Appl.  {\bf 26}, 987--999 (2020)]. In addition, we obtain transformational identities by the method of $q$-difference equation. Moreover, we deduce $U(n+1)$ type generating functions and Ramanujan's integrals involving general Al-Salam--Carlitz polynomials by $q$-difference equation.
\end{abstract}

\maketitle

\section{Introduction}

In this paper, we refer to the general references \cite{GasparRahman} for definitions and notations. Throughout this paper, we suppose that $0<q<1$. For complex numbers $a$, the $q$-shifted factorials are defined by:
$$ (a;q)_n=\left\{
\begin{aligned}
1,\quad\quad\quad\quad\quad\quad\quad \quad n&=0  \\
(1-a)(1-aq)...(1-aq^{n-1}),\quad n&=1,2,3...
\end{aligned}
\right.
$$
and $(a_1,a_2,...,a_m;q)_n=(a_1;q)_n(a_2;q)_n...(a_m;q)_n$, where $m$ is a positive integer and $n$ is a non-negative integer or $\infty$.\par
    The $q$-binomial coefficient is defined by
\begin{equation}
    \begin{bmatrix}
                        n \\
                       k \\
    \end{bmatrix}=\frac{(q;q)_n}{(q;q)_k(q;q)_{n-k}}.
\end{equation}\par
 
The basic (or $q$-) hypergeometric function 
of the variable $z$ and with $\mathfrak{r}$ numerator 
and $\mathfrak{s}$ denominator parameters 
is defined as follows (see, for details, the monographs by 
\mbox{Slater~(\cite{Slater}, Chapter 3)}
and by Srivastava and Karlsson~(\cite{SrivastavaKarlsson}, 
p. 347, Eq. (272)); 
see also \cite{Koekock,HMS-IMAJAM1983-1984a,
HMS-IMAJAM1983-1984b}): 
 
$${}_{\mathfrak r}\Phi_{\mathfrak s}\left[
\begin{array}{rr}
a_1, a_2,\cdots, a_{\mathfrak r};\\
\\
b_1,b_2,\cdots,b_{\mathfrak s};
\end{array}\,
q;z\right]
:=\sum_{n=0}^\infty\Big[(-1)^n \; 
q^{\binom{n}{2}}\Big]^{1+{\mathfrak s}-{\mathfrak r }}
\,\frac{(a_1, a_2,\cdots, a_{\mathfrak r};q)_n}
{(b_1,b_2,\cdots,b_{\mathfrak s};q)_n}
\; \frac{z^n}{(q;q)_n},$$
where $q\neq 0$ when ${\mathfrak r }>{\mathfrak s}+1$. 
We also note that
$${}_{\mathfrak r+1}\Phi_{\mathfrak r}\left[
\begin{array}{rr}
a_1, a_2,\cdots, a_{\mathfrak r+1}\\
\\
b_1,b_2,\cdots,b_{\mathfrak r };
\end{array}\,
q;z\right]
=\sum_{n=0}^\infty \frac{(a_1, a_2,\cdots, a_{\mathfrak r+1};q)_n}
{(b_1,b_2,\cdots,b_{\mathfrak r};q)_n}\;\frac{ z^n}{(q;q)_n}.$$\par
We remark in passing that, in a recently-published 
survey-cum-expository review article, the~so-called $(p,q)$-calculus 
was exposed to be a rather trivial and inconsequential variation of 
the classical $q$-calculus, the additional parameter $p$ being redundant 
or superfluous \mbox{(see, for details,~(\cite{HMS-ISTT2020}, p. 340)).} 
\par
 
The basic (or $q$-) 
series and basic (or $q$-) polynomials, especially
the basic (or $q$-) hypergeometric functions and basic 
(or $q$-) hypergeometric polynomials, are 
applicable particularly in several diverse areas 
[see also~(\cite{SrivastavaKarlsson}, pp. 350--351)]. In particular, the
celebrated Chu-Vandermonde summation theorem and its known $q$-extensions,
which   have already been 
demonstrated to be useful (see, for details, 
\cite{GasparRahman,Srivastava2012,GEAndrews1986,HMS-JC-SA2020}).

    The usual $q$-differential operator, or $q$-derivative, is defined by \cite{14,15,SrivastaAbdlhusein}
\begin{align}
    D_a\{f(a)\}=\frac{f(a)-f(aq)}{a},\quad
    \theta_a\{f(a)\}=\frac{f(aq^{-1})-f(a)}{q^{-1}a}.
\end{align}\par
    The Leibniz rule for $D_a$ and $\theta_a$ are the following identities \cite{14,15,11}
\begin{align}
    D_a^n\{f(a)g(a)\}&=\sum_{k=0}^nq^{k(k-n)}\begin{bmatrix}n \\k \\ \end{bmatrix}D_a^k\{f(a)\}D_a^{n-k}\Bigl\{g\bigl(aq^k\bigr)\Bigr\},\\
     \theta_a^n\{f(a)g(a)\}&=\sum_{k=0}^n\begin{bmatrix}n \\k \\ \end{bmatrix}\theta_a^k\{f(a)\}\theta_a^{n-k}\Bigl\{g\bigl(aq^{-k}\bigr)\Bigr\}.
\end{align} \par
    The following property of $D_q$ is straightforward and important \cite{18}
\begin{align*}
    D_q\left\{\frac{1}{(at;q)_\infty}\right\}=\frac{t}{(at;q)_\infty},\quad
    D_q^k\left\{\frac{1}{(at;q)_\infty}\right\}=\frac{t^k}{(at;q)_\infty},
\end{align*}
\begin{equation}
D_q^k\{a^n\}=\left\{
\begin{aligned}
\frac{(q,q)_n}{(q,q)_{n-k}}a^{n-k},&\quad0\leq k\leq n-1  \\
(q,q)_n,\quad\quad\quad &\quad k=n \\
0,\quad\quad\quad\quad\quad &\quad k\geq n+1.
\end{aligned}
\right.
\end{equation}\par
    The Al-Salam--Carlitz polynomials were introduced by Al-Salam and Carlitz in 1965 \cite[Eqs. (1.11) and (1.15)]{1}
\begin{align}
   \phi_n^{(a)}(x|q)=\sum_{k=0}^n\begin{bmatrix}
                        n \\
                       k \\
    \end{bmatrix}(a;q)_kx^k\quad and\quad
    \psi_n^{(a)}(x|q)=\sum_{k=0}^n\begin{bmatrix}
                        n \\
                       k \\
    \end{bmatrix}q^{k(k-n)}(aq^{1-k};q)_kx^k.
\end{align}\par
    They play important roles in the theory of $q$-orthogonal polynomials. In fact, there are two families of these polynomials: one with continuous orthogonality and another with discrete orthogonality, which given explicitly in the book of Koekoek--Swarttouw--Lesky {\cite[Eqs. (14.24.1) and (14.25.1)]{5}}. For further information about $q$-polynomials, see \cite{cao-niu2016,cao-sri-liu,cao,5,wang2012,wang2013}.\par
    The generalized Al-Salam--Carlitz polynomials \cite[Eq. (4.7)]{17}
\begin{align}
\phi_n^{(a,b,c)}(x,y|q)&=\sum_{k=0}^n\begin{bmatrix}
                        n \\
                       k \\
    \end{bmatrix}\frac{(a,b;q)_k}{(c;q)_k}x^ky^{n-k},\nonumber\\
\psi_n^{(a,b,c)}(x,y|q)&=\sum_{k=0}^n\begin{bmatrix}
                        n \\
                       k \\
    \end{bmatrix}\frac{(-1)^kq^{{k+1\choose2}-nk}(a,b;q)_k}{(c;q)_k}x^ky^{n-k},
\end{align}
 whose generating functions are \cite[Eqs. (4.10) and (4.11))]{17}
\begin{align}
\sum_{n=0}^\infty \phi_n^{(a,b,c)}(x,y|q)\frac{t^n}{(q;q)_n}&=\frac{1}{(yt;q)_\infty}{}_2\phi_1\Biggl[\begin{matrix}
\begin{array}{r}
a,b;\\\\
c;
\end{array}
\end{matrix}q;xt\Biggr],\quad \max\{\abs{yt},\abs{xt}\}<1,\label{9}\\
\sum_{n=0}^\infty \psi_n^{(a,b,c)}(x,y|q)\frac{(-1)^nq^{n\choose 2}t^n}{(q;q)_n}&=(yt;q)_\infty{}_2\phi_1\Biggl[\begin{matrix}
\begin{array}{r}
a,b;\\\\
c;
\end{array}
\end{matrix}q;xt\Biggr],\quad \abs{xt}<1.\label{10}
\end{align}
\par
Chen and Liu \cite{14,15} gave the clever way of parameter augmentation by use of the following two $q$-exponential operators
\begin{align}
   \mathbb{T}(bD_a)=\sum_{n=0}^\infty\frac{(bD_a)^n}{(q;q)_n},\quad
   \mathbb{E}(b\theta_a)=\sum_{n=0}^\infty\frac{q^{n\choose2}(b\theta_a)^n}{(q;q)_n},
\end{align}
which is rich and powerful tool for basic hypergeometric series, especially makes many famous results easily fall into this framework. For further information about $q$-exponential operators, see \cite{14,15,fang2013,fang2014,fang2015,jia2014}.\par
Recently, Srivastava, Arjika and Sherif Kelil \cite{6}  introduced the following homogeneous $q$-difference operator $\widetilde{E}(a,b;D_q)$ by
\begin{equation}\label{Srivastava-Arjika-Kelil }
   \widetilde{E}(a,b;D_q)=\sum_{n=0}^\infty\frac{(-1)^nq^{n\choose2}(a;q)_n}{(q;q)_n}(bD_a)^n.
\end{equation}
The operators \eqref{Srivastava-Arjika-Kelil } have turned out to be suitable for dealing with a generalized Cauchy polynomials $p_n(x,y,a)$
\begin{equation}
  p_n(x,y,a)=\widetilde{E}(a,b;D_q)\{x^n\}.
\end{equation}
For more information about the relations between operators and $q$-polynomials, see \cite{6}.\par
Liu \cite{18,16} deduced several results involving Bailey's $_6\psi_6$, $q$-Mehler formulas for Rogers--Szeg\"{o} polynomials and $q$-integral of Sears' transformation by the following $q$-difference equations.

\begin{pro}[{\cite[Theorems 1 and 2]{16}}]\label{pro1}
Let $f(a,b)$ be a two-variable analytic function in a neighbourhood of $(a,b)=(0,0)\in\mathbb{C}^2$\par
(I) If $f(a,b)$ satisfied the difference equation
\begin{equation}
bf(aq,b)-af(a,bq)=(b-a)f(a,b),
\end{equation}\par
\quad\quad then we have
\begin{equation}
f(a,b)=\mathbb{T}(bD_a)\{f(a,0)\}.
\end{equation}\par
(II) If $f(a,b)$ satisfied the difference equation
\begin{equation}
af(aq,b)-bf(a,bq)=(a-b)f(aq,bq),
\end{equation}\par
\quad\quad then we have
\begin{equation}
f(a,b)=\mathbb{E}(b\theta_a)\{f(a,0)\}.
\end{equation}
\end{pro}
 Arjika \cite{10} continue to consider the following generalized $q$-difference equations.
\begin{pro}[{\cite[Theorem 2.1]{10}}]\label{pro3}
Let $f(a,x,y)$ be a three-variable analytic function in a neighbourhood of $(a,x,y)=(0,0,0)\in\mathbb{C}^3$. If $f(a,x,y)$ can be expanded in terms of $p_n(x,y,a)$ if and only if
\begin{equation}\label{pro3_1}
\quad\quad\quad x[f(a,x,y)-f(a,x,qy)]=y[f(a,qx,qy)-f(a,x,qy)]-ay[f(a,qx,q^2y)-f(a,x,q^2y)].
\end{equation}
\end{pro}
\begin{pro}[{\cite[Theorem 2.2]{10}}]\label{pro2}
Let $f(a,x,y)$ be a three-variable analytic function in a neighbourhood of $(a,x,y)=(0,0,0)\in\mathbb{C}^3$.
If $f(a,x,y)$ satisfied the $q$-difference equation
\begin{equation}\label{pro2_1}
\quad\quad\quad x[f(a,x,y)-f(a,x,qy)]=y[f(a,qx,qy)-f(a,x,qy)]-ay[f(a,qx,q^2y)-f(a,x,q^2y)],
\end{equation}\par
then we have
\begin{equation}
f(a,x,y)=\widetilde{E}(a,b;D_q)\{f(a,x,0)\}.
\end{equation}
\end{pro}
In this paper, our goal is to generalize the results of   Arjika \cite{10}  in section 2. We first construct the following $q$-operators
\begin{align}\label{1.9}
   \mathbb{T}(a,b,c,d,e,yD_x)&=\sum_{n=0}^\infty\frac{(a,b,c;q)_n}{(q,d,e;q)_n}(yD_x)^n,\\
   \mathbb{E}(a,b,c,d,e,y\theta_x)&=\sum_{n=0}^\infty\frac{(-1)^nq^{n\choose2}(a,b,c;q)_n}{(q,d,e;q)_n}(y\theta_x)^n.
\end{align}

We remark that the $q$-operator (\ref{1.9}) is a particular case of  
the homogeneous $q$-difference operator $\mathbb{T}({\bf a},{\bf b},cD_{x})$  
(see \cite{HMS-Sama2020}) by taking  
$${\bf a} =(a,b,c),\quad {\bf b} =(d,e) \qquad \text{and} \qquad c=y.$$

We also built the relations between operators $\mathbb{T}(a,b,c,d,e,yD_x)$, $\mathbb{E}(a,b,c,d,e,y\theta_x)$ and  the new generalized Al-Salam--Carlitz polynomials $\phi_n^{a,b,c\choose d,e}(x,y|q)$, $\psi_n^{a,b,c\choose d,e}(x,y|q),$ respectively,
\begin{align}
\phi_n^{a,b,c\choose d,e}(x,y|q)\triangleq\mathbb{T}(a,b,c,d,e,yD_x)\{x^n\}&=\sum_{k=0}^n\begin{bmatrix}
                        n \\
                       k \\
    \end{bmatrix}\frac{(a,b,c;q)_k}{(d,e;q)_k}x^{n-k}y^k,\label{23}\\
\psi_n^{a,b,c\choose d,e}(x,y|q)\triangleq\mathbb{E}(a,b,c,d,e,y\theta_x)\{x^n\}&=\sum_{k=0}^n\begin{bmatrix}
                        n \\
                       k \\
    \end{bmatrix}\frac{(-1)^kq^{k(k-n)}(a,b,c;q)_k}{(d,e;q)_k}x^{n-k}y^k.\label{24}
\end{align} 

The paper is organized as follows: In section \ref{section2}, we state two theorems and give the proofs. In section \ref{section3}, we gain generalize generating functions for new generalized Al-Salam--Carlitz polynomials by using the method of $q$-difference equations perspectively. In section \ref{section4}, we obtain a transformational identities involving generating functions for generalized Al-Salam--Carlitz polynomials by $q$-difference equations. In section \ref{section5}, we deduce $U(n+1)$ type generating functions for generalized Al-Salam--Carlitz polynomials by $q$-difference equation. In section \ref{section6}, we deduce generalizations of Ramanujan's integrals.

\section{Main results and proofs}\label{section2}

In this section, we give the following two theorems.
\begin{thm}\label{thm1}
Let $f(a,b,c,d,e,x,y)$ be a seven-variable analytic function in a neighbourhood of $(a,b,c,d,e,x,y)=(0,0,0,0,0,0,0)\in\mathbb{C}^7$.\par
(I) If $f(a,b,c,d,e,x,y)$ can be expanded in terms of $\phi_n^{a,b,c\choose d,e}(x,y|q)$ if and only if
\begin{align}\label{thm1_1}
&\quad\quad x\{f(a,b,c,d,e,x,y)-f(a,b,c,d,e,x,yq)-(d+e)q^{-1}[f(a,b,c,d,e,x,yq)-f(a,b,c,d,e,x,yq^2)]\nonumber\\
&\quad\quad\quad +deq^{-2}[f(a,b,c,d,e,x,yq^2)-f(a,b,c,d,e,x,yq^3)]\}\nonumber\\
&=y\{[f(a,b,c,d,e,x,y)-f(a,b,c,d,e,xq,y)]-(a+b+c)[f(a,b,c,d,e,x,yq)-f(a,b,c,d,e,xq,yq)]\nonumber\\
&\quad\quad\quad +(ab+ac+bc)[f(a,b,c,d,e,x,yq^2)-f(a,b,c,d,e,xq,yq^2)]\nonumber\\
&\quad\quad\quad\quad  -abc[f(a,b,c,d,e,x,yq^3)-f(a,b,c,d,e,xq,yq^3)]\}.
\end{align}\par
(II) If $f(a,b,c,d,e,x,y)$ can be expanded in terms of $\psi_n^{a,b,c\choose d,e}(x,y|q)$ if and only if
\begin{align}\label{thm1_2}
&\quad\quad x\{f(a,b,c,d,e,xq,y)-f(a,b,c,d,e,xq,yq)-(d+e)q^{-1}[f(a,b,c,d,e,xq,yq)-f(a,b,c,d,e,xq,yq^2)]\nonumber\\
&\quad\quad\quad +deq^{-2}[f(a,b,c,d,e,xq,yq^2)-f(a,b,c,d,e,xq,yq^3)]\}\nonumber\\
&=y\{[f(a,b,c,d,e,xq,yq)-f(a,b,c,d,e,x,yq)]-(a+b+c)[f(a,b,c,d,e,xq,yq^2)-f(a,b,c,d,e,x,yq^2)]\nonumber\\
&\quad\quad\quad +(ab+ac+bc)[f(a,b,c,d,e,xq,yq^3)-f(a,b,c,d,e,x,yq^3)]\nonumber\\
&\quad\quad\quad\quad  -abc[f(a,b,c,d,e,xq,yq^4)-f(a,b,c,d,e,x,yq^4)]\}.
\end{align}
\end{thm}

\begin{rem}
For $c$=$d$=$e$=$0$, and $b$$\rightarrow$ $\frac{1}{b}$,  $y\rightarrow yb$, $b\rightarrow 0$, then equation \eqref{thm1_1}  reduces  \eqref{pro3_1}.
\end{rem}

\begin{thm}\label{thm2}
Let $f(a,b,c,d,e,x,y)$ be a seven-variable analytic function in a neighbourhood of $(a,b,c,d,e,x,y)=(0,0,0,0,0,0,0)\in\mathbb{C}^7$.

(I) If $f(a,b,c,d,e,x,y)$ satisfied the difference equation
\begin{align}\label{thm2_1}
&\quad\quad x\{f(a,b,c,d,e,x,y)-f(a,b,c,d,e,x,yq)-(d+e)q^{-1}[f(a,b,c,d,e,x,yq)-f(a,b,c,d,e,x,yq^2)]\nonumber\\
&\quad\quad\quad +deq^{-2}[f(a,b,c,d,e,x,yq^2)-f(a,b,c,d,e,x,yq^3)]\}\nonumber\\
&=y\{[f(a,b,c,d,e,x,y)-f(a,b,c,d,e,xq,y)]-(a+b+c)[f(a,b,c,d,e,x,yq)-f(a,b,c,d,e,xq,yq)]\nonumber\\
&\quad\quad\quad +(ab+ac+bc)[f(a,b,c,d,e,x,yq^2)-f(a,b,c,d,e,xq,yq^2)]\nonumber\\
&\quad\quad\quad\quad  -abc[f(a,b,c,d,e,x,yq^3)-f(a,b,c,d,e,xq,yq^3)]\},
\end{align}
then we have
\begin{align}\label{thm2_1.2}
f(a,b,c,d,e,x,y)=\mathbb{T}(a,b,c,d,e,yD_x)\{f(a,b,c,d,e,x,0)\}.
\end{align}
(II)  If $f(a,b,c,d,e,x,y)$ satisfied the difference equation
\begin{align}\label{thm2_2}
&\quad\quad x\{f(a,b,c,d,e,xq,y)-f(a,b,c,d,e,xq,yq)-(d+e)q^{-1}[f(a,b,c,d,e,xq,yq)-f(a,b,c,d,e,xq,yq^2)]\nonumber\\
&\quad\quad\quad +deq^{-2}[f(a,b,c,d,e,xq,yq^2)-f(a,b,c,d,e,xq,yq^3)]\}\nonumber\\
&=y\{[f(a,b,c,d,e,xq,yq)-f(a,b,c,d,e,x,yq)]-(a+b+c)[f(a,b,c,d,e,xq,yq^2)-f(a,b,c,d,e,x,yq^2)]\nonumber\\
&\quad\quad\quad +(ab+ac+bc)[f(a,b,c,d,e,xq,yq^3)-f(a,b,c,d,e,x,yq^3)]\nonumber\\
&\quad\quad\quad\quad  -abc[f(a,b,c,d,e,xq,yq^4)-f(a,b,c,d,e,x,yq^4)]\},
\end{align}
then we have
\begin{align}
f(a,b,c,d,e,x,y)=\mathbb{E}(a,b,c,d,e,y\theta_x)\{f(a,b,c,d,e,x,0)\}.
\end{align}
\end{thm}

\begin{rem}
For $c$=$d$=$e$=$0$, and $b$$\rightarrow$ $\frac{1}{b}$,  $y\rightarrow yb$, $b\rightarrow 0$, then equation \eqref{thm2_1} reduces to \eqref{pro2_1}.
\end{rem}

To determine if a given function is an analytic function in several complex variables, we often use the following Hartogs's theorem. For more information, please refer to \cite{19,7}.

\begin{lem}[{{\cite[Hartogs's theorem]{9}}}]
If a complex-valued function is holomorphic (analytic) in each variable separately in an open domain $D \in\mathbb{C}^n$, then it is holomorphic (analytic) in $D$.
\end{lem}

In order to prove Theorem \ref{thm1}, we need the following fundamental property of several complex variables.
\begin{lem}[{\cite[Proposition 1]{3}}]
If $f(x_1,x_2,...,x_k)$ is analytic at the origin $(0,0,...,0)\in\mathbb{C}^k$, then, $f$ can be expanded in an absolutely convergent power series,
\begin{align*}
f(x_1,x_2,...,x_k)=\sum_{n_1,n_2,...,n_k=0}^\infty\alpha_{n_1,n_2,...,n_k}x_1^{n_1}x_2^{n_2}...x_k^{n_k}.
\end{align*}
\end{lem}

\begin{proof}[Proof of Theorem \ref{thm1}]
(I) From the Hartogs's theorem and the theory of several complex variables, we assume that
\begin{align}\label{27}
f(a,b,c,d,e,x,y)=\sum_{k=0}^\infty A_k(a,b,c,d,e,x)y^k.
\end{align}\par
On one hand, substituting \eqref{27} into \eqref{thm1_1} yields
\begin{multline}
 x\sum_{k=0}^\infty[1-q^k-(d+e)q^{k-1}+(d+e)q^{2k-1}+deq^{2k-2}-deq^{3k-2}]A_k(a,b,c,d,e,x)y^k\\
=\sum_{k=0}^\infty[1-(a+b+c)q^k+(ab+bc+ac)q^{2k}-abcq^{3k}][A_k(a,b,c,d,e,x)-A_k(a,b,c,d,e,xq)]y^{k+1},
\end{multline}
which is equal to
\begin{multline}\label{A-n rel-eq}
 x\sum_{k=0}^\infty(1-q^k)(1-dq^{k-1})(1-eq^{k-1})A_k(a,b,c,d,e,x)y^k\\
=\sum_{k=0}^\infty(1-aq^{k})(1-bq^{k})(1-cq^{k})[A_k(a,b,c,d,e,x)-A_k(a,b,c,d,e,xq)]y^{k+1}.
\end{multline}\par
Equating coefficients of $y^k$ on both sides of equation \eqref{A-n rel-eq}, we have
\begin{multline}
 x(1-q^k)(1-dq^{k-1})(1-eq^{k-1})A_k(a,b,c,d,e,x)\\
=(1-aq^{k-1})(1-bq^{k-1})(1-cq^{k-1})[A_{k-1}(a,b,c,d,e,x)-A_{k-1}(a,b,c,d,e,xq)],
\end{multline}
which is equivalent to
\begin{align*}
A_k(a,b,c,d,e,x)&=\frac{(1-aq^{k-1})(1-bq^{k-1})(1-cq^{k-1})}{(1-q^k)(1-dq^{k-1})(1-eq^{k-1})}\cdot\frac{A_{k-1}(a,b,c,d,e,x)-A_{k-1}(a,b,c,d,e,xq)}{x} \\
&=\frac{(1-aq^{k-1})(1-bq^{k-1})(1-cq^{k-1})}{(1-q^k)(1-dq^{k-1})(1-eq^{k-1})}\cdot D_x\{A_{k-1}(a,b,c,d,e,x)\}.
\end{align*}\par
By iteration, we gain
\begin{align*}
A_k(a,b,c,d,e,x)=\frac{(a,b,c;q)_k}{(q,d,e;q)_k}\cdot D_x^k\{A_0(a,b,c,d,e,x)\}.
\end{align*}\par
Letting $\displaystyle f(a,b,c,d,e,x,0)=A_0(a,b,c,d,e,x)=\sum_{n=0}^\infty\mu_nx^n$ yields
\begin{align}\label{28}
A_k(a,b,c,d,e,x)=\frac{(a,b,c;q)_k}{(q,d,e;q)_k}\cdot\sum_{n=0}^\infty\mu_n\frac{(q;q)_n}{(q;q)_{n-k}}x^{n-k},
\end{align}
we have
\begin{align*}
f(a,b,c,d,e,x,y)&=\sum_{k=0}^\infty\frac{(a,b,c;q)_k}{(q,d,e;q)_k}\sum_{n=0}^\infty\mu_n\frac{(q;q)_n}{(q;q)_{n-k}}x^{n-k}y^k\\
&=\sum_{n=0}^\infty\mu_n\sum_{k=0}^\infty\begin{bmatrix}
                        n \\
                       k \\
    \end{bmatrix}\frac{(a,b,c;q)_k}{(d,e;q)_k}x^{n-k}y^k\\
&=\sum_{n=0}^\infty\mu_n\phi_n^{a,b,c\choose d,e}(x,y|q).
\end{align*}
On the other hand, if $f(a,b,c,d,e,x,y)$ can be expanded in terms of $\phi_n^{a,b,c\choose d,e}(x,y|q)$, we verify that $f(a,b,c,d,e,x,y)$ satisfies \eqref{thm1_1}. Similarly, we deduce (II). The proof of Theorem \ref{thm1} is complete.
\end{proof}

\begin{proof}[Proof of Theorem \ref{thm2}]
From the theory of several complex variables, we begin to solve the $q$-difference. First we may assume that
\begin{align}\label{33}
f(a,b,c,d,e,x,y)=\sum_{k=0}^\infty A_k(a,b,c,d,e,x)y^k.
\end{align}\par
Substituting this equation into \eqref{33} and compare coefficients of $y^k\,\,(k\geq1)$, we readily find that
\begin{multline}
 x(1-q^k)(1-dq^{k-1})(1-eq^{k-1})A_k(a,b,c,d,e,x)\\
=(1-aq^{k-1})(1-bq^{k-1})(1-cq^{k-1})[A_{k-1}(a,b,c,d,e,x)-A_{k-1}(a,b,c,d,e,xq)],
\end{multline}
which equals
\begin{align*}
A_k(a,b,c,d,e,x)&=\frac{(1-aq^{k-1})(1-bq^{k-1})(1-cq^{k-1})}{(1-q^k)(1-dq^{k-1})(1-eq^{k-1})}\cdot\frac{A_{k-1}(a,b,c,d,e,x)-A_{k-1}(a,b,c,d,e,xq)}{x} \\
&=\frac{(1-aq^{k-1})(1-bq^{k-1})(1-cq^{k-1})}{(1-q^k)(1-dq^{k-1})(1-eq^{k-1})}\cdot D_x\{A_{k-1}(a,b,c,d,e,x)\}.
\end{align*}\par
By iteration, we gain
\begin{align}\label{34}
A_k(a,b,c,d,e,x)=\frac{(a,b,c;q)_k}{(q,d,e;q)_k}\cdot D_x^k\{A_0(a,b,c,d,e,x)\}.
\end{align}\par
Now we return to calculate $A_0(a,b,c,d,e,x)$. Just taking $y=0$ in equation \eqref{33}, we immediately obtain $A_0(a,b,c,d,e,x)=f(a,b,c,d,e,x,0)$. Substituting \eqref{34} into \eqref{33}, we achieve \eqref{thm2_1.2} directly. The proof of Theorem \ref{thm2} is complete.
\end{proof}


\section{Generating functions for new generalized Al-Salam--Carlitz polynomials}\label{section3}
In this section we generalized generating functions for the new generalized Al-Salam--Carlitz polynomials by the method of $q$-difference equations.\par
We start with the following lemmas.
\begin{lem}[\cite{13}]
The Cauchy polynomials as given below
\begin{align}\label{lem_3}
p_n(x,y)=(x-y)(x-qy)...(x-q^{n-1}y)=(y/x;q)_nx^n
\end{align}
together with the following Srivastava-Agarwal 
type generating function 
(see also \cite{Cao-Srivastava2013}):  
\begin{equation}
\label{Srivas}
\sum_{n=0}^\infty  
p_n (x,y)\;\frac{(\lambda;q)_n\,t^n}{(q;q)_n}
={}_{2}\Phi_1\left[
\begin{array}{rr}
\lambda,\frac{y}{x};\\
\\
0;
\end{array}\,
q;  xt\right].
\end{equation}
\end{lem}
\begin{lem}[\cite{13}]\label{lem_4}
Suppose that $max\{|xt|,|yt|\}<1$, we have
\begin{align}
\sum_{n=0}^\infty p_n(x,y)\frac{t^n}{(q;q)_n}=\frac{(yt;q)_\infty}{(xt;q)_\infty}.\label{putt}
\end{align}
\end{lem}

Based upon the $q$-binomial theorem or   the homogeneous version of the Cauchy identity
(\ref{putt}) and Heine's
transformations, Srivastava {et al.} \cite{HMS-C-W2020} established 
a set of two presumably new theta-function identities 
(see, for details, \cite{HMS-C-W2020}).
\begin{lem}[{\cite[Theorem 5]{13}}]
Suppose that $\max\{\abs{act},\abs{adt},\abs{bct},\abs{bdt}\}<1$, we have
\begin{align}\label{lem_5}
\sum_{n=0}^\infty h_n(a,b|q)h_n(c,d|q)\frac{t^n}{(q;q)_n}=\frac{(abcdt^2;q)_\infty}{(act,adt,bct,bdt;q)_\infty}.
\end{align}
\end{lem}
\begin{thm}\label{thm3}
Suppose that $max\{|xt|,|yt|\}<1$, we have
\begin{align}\label{thm3_1}
\sum_{n=0}^\infty \phi_n^{a,b,c\choose d,e}(x,y|q)\frac{t^n}{(q;q)_n}=\frac{1}{(xt;q)_\infty}{}_3\phi_2\Biggl[\begin{matrix}
\begin{array}{r}
a,b,c;\\\\
d,e;
\end{array}
\end{matrix}q;yt\Biggr],
\end{align}
\begin{align}\label{thm3_2}
\sum_{n=0}^\infty \psi_n^{a,b,c\choose d,e}(x,y|q)\frac{t^n}{(q;q)_n}=(xt;q)_\infty{}_3\phi_3\Biggl[\begin{matrix}
\begin{array}{r}
a,b,c;\\\\
0,d,e;
\end{array}
\end{matrix}q;-yt\Biggr].
\end{align}
\end{thm}
\begin{rem}
For $c=e=0$ in Theorem \ref{thm3}, equations \eqref{thm3_1} and \eqref{thm3_2} reduce to equations \eqref{9} and \eqref{10}, respectively.
\end{rem}
\begin{proof}[Proof of Theorem \ref{thm3}]
By denoting the right-hand side of equation \eqref{thm3_1} by $f(a,b,c,d,e,x,y)$,we can verify that $f(a,b,c,d,e,x,y)$ satisfies \eqref{thm1_1}. So, we have
\begin{align*}
f(a,b,c,d,e,x,y)=\sum_{k=0}^\infty\mu_n\phi_n^{a,b,c\choose d,e}(x,y|q),
\end{align*}
and
\begin{align*}
f(a,b,c,d,e,x,0)=\sum_{k=0}^\infty\mu_nx^n=\frac{1}{(xt;q)_\infty}=\sum_{n=0}^\infty\frac{t^n}{(q;q)_n}x^n.
\end{align*}
So, $f(a,b,c,d,e,x,y)$ is equal to
\begin{align*}
f(a,b,c,d,e,x,y)=\sum_{k=0}^\infty\frac{t^n}{(q;q)_n}\phi_n^{a,b,c\choose d,e}(x,y|q),
\end{align*}
equal to the right-hand side of equation \eqref{thm3_1}. \par
Similarly, by denoting the right-hand side of equation \eqref{thm3_2} by $f(a,b,c,d,e,x,y)$, we can verify that $f(a,b,c,d,e,x,y)$ satisfies \eqref{thm1_2}. So, we can use the same way to achieve the equation \eqref{thm3_2}. The proof of Theorem \ref{thm3} is complete.
\end{proof}

\begin{thm}\label{thm4}
Suppose that $\max\{\abs{xt},\abs{yt}\}<1$, we have
\begin{align}\label{thm4_1}
\sum_{n=0}^\infty \phi_n^{a,b,c\choose d,e}(x,y|q)\frac{p_n(s,t)}{(q;q)_n}=\frac{(xs;q)_\infty}{(xt;q)_\infty}{}_4\phi_3\Biggl[\begin{matrix}
\begin{array}{r}
a,b,c,s/t;\\\\
d,e,xs;
\end{array}
\end{matrix}q;yt\Biggr].
\end{align}
\end{thm}
\begin{cor}
Suppose that $\abs{yt}<1$, we have
\begin{align}\label{cor1_1}
\sum_{n=0}^\infty \phi_n^{a,b,c\choose d,e}(x,y|q)\frac{(-1)^nq^{n\choose2}t^n}{(q;q)_n}=(xt;q)_\infty{}_3\phi_3\Biggl[\begin{matrix}
\begin{array}{r}
a,b,c;\\\\
d,e,xt;
\end{array}
\end{matrix}q;yt\Biggr].
\end{align}
\end{cor}
\begin{rem}
For $t=0$, in Theorem \ref{thm4}, equation \eqref{thm4_1} reduces to \eqref{cor1_1}. For $s=0$ in Theorem \ref{thm4}, equation \eqref{thm4_1} reduces to \eqref{thm3_1}, respectively.
\end{rem}

\begin{proof}[Proof of Theorem \ref{thm4}]
By denoting the right-hand side of equation \eqref{thm4_1} by $f(a,b,c,d,e,x,y)$, we can verify that $f(a,b,c,d,e,x,y)$ satisfies \eqref{thm1_1}. So, we have
\begin{align}
f(a,b,c,d,e,x,y)=\sum_{k=0}^\infty\mu_n\phi_n^{a,b,c\choose d,e}(x,y|q),
\end{align}
and
\begin{align*}
f(a,b,c,d,e,x,0)=\sum_{k=0}^\infty\mu_nx^n=\frac{(xs;q)_\infty}{(xt;q)_\infty}=\sum_{n=0}^\infty\frac{p_n(t,s)}{(q;q)_n}x^n.
\end{align*}
So, $f(a,b,c,d,e,x,y)$ is equal to the right-hand side of equation \eqref{thm4_1}. The proof of Theorem \ref{thm4} is complete.
\end{proof}

\begin{thm}\label{thm5}
For $k\in\mathbb{N}$ and $max\{|xt|,|yt|\}<1$, we have
\begin{align}\label{thm5_1}
\sum_{n=0}^\infty \phi_{n+k}^{a,b,c\choose d,e}(x,y|q)\frac{t^n}{(q;q)_n}=\frac{x^k}{(xt;q)_\infty}\sum_{n=0}^\infty\frac{(a,b,c;q)_n}{(q,d,e;q)_n}(yt)^n\sum_{j=0}^n\begin{bmatrix}
                        n \\
                       k \\
    \end{bmatrix}\frac{(-1)^jq^{nj-{j\choose2}}(q^{-k},xt;q)_j}{(xt)^j}.
\end{align}
\end{thm}
\begin{rem}
For $k=0$ in Theorem \ref{thm5}, equation \eqref{thm5_1} reduces to \eqref{thm3_1}.
\end{rem}
\begin{proof}[Proof of Theorem \ref{thm5}]
Denoting the right-hand side of equation \eqref{thm5_1} equivalently by
\begin{align}\label{44}
f(a,b,c,d,e,x,y)=\frac{x^k}{(xt;q)_\infty}\sum_{n=0}^\infty\frac{(a,b,c;q)_n}{(q,d,e;q)_n}(yt)^n\sum_{j=0}^n\begin{bmatrix}
                        n \\
                       k \\
    \end{bmatrix}\frac{(-1)^jq^{nj-{j\choose2}}(q^{-k},xt;q)_j}{(xt)^j},
\end{align}
and it is easy to check that \eqref{44} satisfies \eqref{thm1_1}, so we have
\begin{align}\label{45}
f(a,b,c,d,e,x,y)=\sum_{k=0}^\infty\mu_n\phi_n^{a,b,c\choose d,e}(x,y|q).
\end{align}
Setting $y=0$ in \eqref{45}, it becomes
\begin{align*}
f(a,b,c,d,e,x,0)=\sum_{k=0}^\infty\mu_nx^n=\frac{x^n}{(xt;q)_\infty}=\sum_{n=0}^\infty x^k\frac{(xt)^n}{(q;q)_n}=\sum_{n=0}^\infty x^{n+k}\frac{t^n}{(q;q)_n}=\sum_{n=k}^\infty x^n\frac{t^{n-k}}{(q;q)_{n-k}}.
\end{align*}
Hence
\begin{align*}
f(a,b,c,d,e,x,y)=\sum_{n=k}^\infty\phi_n^{a,b,c\choose d,e}(x,y|q)\frac{t^{n-k}}{(q;q)_{n-k}}=\sum_{n=0}^\infty \phi_{n+k}^{a,b,c\choose d,e}(x,y|q)\frac{t^n}{(q;q)_n}.
\end{align*}\par
The proof of Theorem \ref{thm5} is complete.
\end{proof}
\begin{thm}\label{thm6}
We have
\begin{multline}\label{thm6_1}
 \sum_{n=0}^\infty \phi_n^{a_1,b_1,c_1\choose d_1,e_1}(x_1,y_1|q)\phi_n^{a_2,b_2,c_2\choose d_2,e_2}(x_2,y_2|q)\frac{t^n}{(q;q)_n}=\frac{1}{(x_1x_2t;q)_\infty}\sum_{n=0}^\infty\frac{(a_2,b_2,c_2;q)_n(x_1y_2t)^n}{(q,d_2,e_2;q)_n}\\
\times\sum_{j=0}^\infty\frac{(q^{n-j+1},x_1x_2t,a_1,b_1,c_1;q)_j(\frac{y_1}{x_1})^j}{(q,d_1,e_1;q)_j}
{}_3\phi_2\Biggl[\begin{matrix}
\begin{array}{r}
a_1q^j,b_1q^j,c_1q^j;\\\\
d_1q^j,e_1q^j;
\end{array}
\end{matrix}q;x_2y_1t\Biggr].
\end{multline}
\end{thm}
\begin{rem}
For $a_1$=$b_1$=$c_1$=$d_1$=$e_1$=$a_2$=$b_2$=$c_2$=$d_2$=$e_2$=$0$ in Theorem \ref{thm6}, equation \eqref{thm6_1} reduces to \eqref{lem_5}.
\end{rem}
\begin{proof}[Proof of Theorem \ref{thm6}]
Denoting the right-hand side of equation \eqref{thm6_1} by $H(a_1,b_1,c_1,d_1,e_1,x_1,y_1)$, we have
\begin{multline}\label{47}
\quad H(a_1,b_1,c_1,d_1,e_1,x_1,y_1)=\frac{1}{(x_1x_2t;q)_\infty}\sum_{n=0}^\infty\frac{(a_2,b_2,c_2;q)_n(x_1y_2t)^n}{(q,d_2,e_2;q)_n}\\
\times\sum_{j=0}^\infty\frac{(q^{n-j+1},x_1x_2t,a_1,b_1,c_1;q)_j(\frac{y_1}{x_1})^j}{(q,d_1,e_1;q)_j}
{}_3\phi_2\Biggl[\begin{matrix}
\begin{array}{r}
a_1q^j,b_1q^j,c_1q^j;\\\\
d_1q^j,e_1q^j;
\end{array}
\end{matrix}q;x_2y_1t\Biggr].
\end{multline}\par
Because equation \eqref{47} satisfies \eqref{thm2_1}, we have
\begin{align*}
H(a_1,b_1,c_1,d_1,e_1,x_1,y_1)&=\mathbb{T}(a_1,b_1,c_1,d_1,e_1,y_1D_{x_1})\{H(a_1,b_1,c_1,d_1,e_1,x_1,0)\}\nonumber\\
&=\mathbb{T}(a_1,b_1,c_1,d_1,e_1,y_1D_{x_1})\left\{\frac{1}{(x_1x_2t;q)_\infty}\sum_{n=0}^\infty\frac{(a_2,b_2,c_2;q)_n(x_1y_2t)^n}{(q,d_2,e_2;q)_n}\right\}\\
&=\mathbb{T}(a_1,b_1,c_1,d_1,e_1,y_1D_{x_1})\left\{\sum_{n=0}^\infty\phi_n^{a_2,b_2,c_2\choose d_2,e_2}(x_2,y_2|q)\frac{(x_1t)^n}{(q;q)_n}\right\}\\
&=\sum_{n=0}^\infty\phi_n^{a_2,b_2,c_2\choose d_2,e_2}(x_2,y_2|q)\frac{t^n}{(q;q)_n}\mathbb{T}(a_1,b_1,c_1,d_1,e_1,y_1D_{x_1})\{x_1^n\}\\
&=\sum_{n=0}^\infty\phi_n^{a_1,b_1,c_1\choose d_1,e_1}(x_1,y_1|q) \phi_n^{a_2,b_2,c_2\choose d_2,e_2}(x_2,y_2|q)\frac{t^n}{(q;q)_n}.
\end{align*}\par
The proof of Theorem \ref{thm6} is complete.
\end{proof}

\section{Transformational identities from $q$-difference equations }\label{section4}
Liu \cite{17} gave some important transformational identities by the method of $q$-difference operator. For more details, please refer to \cite{5,17,20}.\par
In this section we deduce the following transformational identities involving generating functions for new generalized Al-Salam--Carlitz polynomials by the method of $q$-difference equation.
\begin{thm}\label{thm7}
Let $A(k)$ and $B(k)$ satisfy
\begin{align}\label{thm7_1}
\sum_{k=0}^\infty A(k)x^k=\sum_{k=0}^\infty B(k)\frac{(xtq^k;q)_\infty}{(xq^k;q)_\infty},
\end{align}
and we have
\begin{align}\label{thm7_2}
\sum_{k=0}^\infty A(k)\phi_k^{a,b,c\choose d,e}(x,y|q)=\sum_{k=0}^\infty B(k)\frac{(xtq^k;q)_\infty}{(xq^k;q)_\infty}{}_4\phi_3\Biggl[\begin{matrix}
\begin{array}{r}
a,b,c,1/t;\\\\
d,e,xtq^k;
\end{array}
\end{matrix}q;yq^k\Biggr],
\end{align}
supposing that equations \eqref{thm7_1} and \eqref{thm7_2} are convergent.
\end{thm}
\begin{cor}\label{cor2}
Suppose that $|r|,|x|,|xt|<1$, we have
\begin{align}\label{cor2_1}
\sum_{k=0}^\infty\phi_k^{a,b,c\choose d,e}(x,y|q)\frac{(t,s;q)_k}{(q,r;q)_k}=\frac{(xt,s;q)_\infty}{(x,r;q)_\infty}\sum_{k=0}^\infty\frac{(r/s,x;q)_ks^k}{(q,xt;q)_k}{}_4\phi_3\Biggl[\begin{matrix}
\begin{array}{r}
a,b,c,1/t;\\\\
d,e,xtq^k;
\end{array}
\end{matrix}q;yq^k\Biggr].
\end{align}
\end{cor}
\begin{rem}
Setting $A(k)$ and $B(k)$ in Theorem \ref{thm7} by \eqref{53} given below, equation \eqref{thm7_2} reduces to \eqref{cor2_1}. For $y=0$ in \eqref{cor2_1}, equation \eqref{thm7_2}  reduces to \eqref{52} below.
\end{rem}
\begin{proof}[Proof of Theorem \ref{thm7}]
Denoting the right-hand side of equation \eqref{thm7_2} equivalently by $f(a,b,c,d,e,x,y)$, and we can check that $f(a,b,c,d,e,x,y)$ satisfied \eqref{thm1_1}, so we have
\begin{align}\label{51}
f(a,b,c,d,e,x,y)=\sum_{k=0}^\infty\mu_n\phi_n^{a,b,c\choose d,e}(x,y|q)
\end{align}\\
Setting $y=0$ in \eqref{51}, it becomes
\begin{align*}
f(a,b,c,d,e,x,0)&=\sum_{k=0}^\infty\mu_nx^n=\sum_{k=0}^\infty B(k)\frac{(xtq^k;q)_\infty}{(xq^k;q)_\infty}\quad by\,\,(48)\\
&=\sum_{k=0}^\infty A(k)x^k.
\end{align*}
Hence
\begin{align*}
f(a,b,c,d,e,x,y)=\sum_{k=0}^\infty A(k)\phi_k^{a,b,c\choose d,e}(x,y|q).
\end{align*}\par
The proof of Theorem \ref{thm7} is complete.
\end{proof}
\begin{proof}[Proof of Corollary \ref{cor2}]
Using the Heine's $q$-Euler transformations \cite[Eq.(1.4.1)]{1}
\begin{align}\label{52}
{}_2\phi_1\Biggl[\begin{matrix}\begin{array}{r}t,s;\\\\r;\end{array}\end{matrix}q;x\Biggr]=\frac{(s,xt;q)_\infty}{(r,x;q)_\infty}
{}_2\phi_1\Biggl[\begin{matrix}\begin{array}{r}r/s,x;\\\\xt;\end{array}\end{matrix}q;s\Biggr],
\end{align}
formula \eqref{thm7_1} is valid for the case
\begin{align}\label{53}
A(k)=\sum_{k=0}^\infty\frac{(t,s;q)_k}{(q,r)_k}\quad and \quad B(k)=\frac{(s;q)_\infty}{(r;q)_\infty}\sum_{k=0}^\infty\frac{(r/s;q)_k}{(q;q)_k}s^k.
\end{align}\par
Using equation \eqref{thm7_2}, we can deduce the Corollary \ref{cor2}.
\end{proof}

\section{$U(n+1)$ type generating functions for generalized Al-Salam-Carlitz polynomials }\label{section5}
    Multiple basic hypergeometric series associated to the unitary $U(n+1)$  group have been investigated by various authors, see \cite{8,12}. In \cite{8}, Milne initiated theory and application of the $U(n+1)$ generalization of the classical Bailey transform and Bailey lemma, which involves the following nonterminating $U(n+1)$  generalizations of the $q$-binomial theorem.
\begin{pro}[{\cite[Theorem 5.42]{18}}]
Let $b,z$ and $x_1,...,x_n$ be indeterminate, and let $n\geq 1$. Suppose that none of the denominators in the following identity vanishes, and that $0<|q|<1$ and $|z|<|x_1,...,x_n||x_m|^{-n}|q|^{(n-1)/2}$, for $m=1,2,...,n$. Then
\begin{multline}\label{lem6_1}
\sum_{\mbox{\tiny$\begin{array}{c} y_n \geq 0\\ k=1,2,...,n\end{array}$}}\biggl\{\prod_{1\leq r<s\leq n}\biggl[\frac{1-(x_r/x_sq^{y_r-y_s})}{1-(x_r/x_s)}\biggl]\prod_{r,s=1}^n(q\frac{x_r}{x_s};q)^{-1}_{y_r}\prod_{i=1}^n(x_i)^{ny_i-(y_1+...+y_n)}(-1)^{(n-1)(y_1+...+y_n)}\\
\times q^{y_2+2y_3+...+(n-1)y_n+(n-1)[{y_1\choose 2}+...+{y_n\choose 2}]-e_2(y_1,...,y_n)}(b;q)_{y_1+...+y_n}z^{y_1+...+y_n}=\frac{(bz;q)_\infty}{(z;q)_\infty},
\end{multline}
where $e_2(y_1,...,y_n)$ is the second elementary symmetric function of $\{y_1,...,y_n\}$.
\end{pro}

In this section, we deduce $U(n+1)$ type generating functions for generalized Al-Salam-Carlitz polynomials by the methodn of $q$-difference equation.
\begin{thm}\label{thm8}
Let $b,z$ and $x_1,...,x_n$ be indeterminate, and let $n\geq 1$. Suppose that none of the denominators in the following identity vanishes, and that $0<|q|<1$ and $|z|<|x_1,...,x_n||x_m|^{-n}|q|^{(n-1)/2}$, for $m=1,2,...,n$. Then
\begin{multline}\label{thm8_1}
\sum_{\mbox{\tiny$\begin{array}{c} y_n \geq 0\\ k=1,2,...,n\end{array}$}}\biggl\{\prod_{1\leq r<s\leq n}\biggl[\frac{1-(x_r/x_sq^{y_r-y_s})}{1-(x_r/x_s)}\biggl]\prod_{r,s=1}^n(q\frac{x_r}{x_s};q)^{-1}_{y_r}\prod_{i=1}^n(x_i)^{ny_i-(y_1+...+y_n)}(-1)^{(n-1)(y_1+...+y_n)}\\
\times q^{y_2+2y_3+...+(n-1)y_n+(n-1)[{y_1\choose 2}+...+{y_n\choose 2}]-e_2(y_1,...,y_n)}\phi_{y_1+...+y_n}^{r,s,t\choose u,v}(z,y|q)(b;q)_{y_1+...+y_n}=\frac{(bz;q)_\infty}{(z;q)_\infty}{}_4\phi_3\Biggl[\begin{matrix}
\begin{array}{r}
r,s,t,b;\\\\
u,v,bz;
\end{array}
\end{matrix}q;y\Biggr],
\end{multline}
where $e_2(y_1,...,y_n)$ is the second elementary symmetric function of $\{y_1,...,y_n\}$.
\end{thm}
\begin{rem}
For $y=0$, in Theorem \ref{thm8}, equation \eqref{thm8_1} reduces to \eqref{lem6_1}.
\end{rem}
\begin{proof}[Proof of Theorem \ref{thm8}]
Denoting the right-hand side of equation \eqref{thm8_1} equivalently by $f(r,s,t,u,v,z,y)$, and we can check that $f(r,s,t,u,v,z,y)$ satisfied \eqref{thm2_1}), so we have
\begin{align*}
f(r,s,t,u,v,z,y)&=\mathbb{T}(r,s,t,u,v,yD_z)\{f(r,s,t,u,v,z,0)\}\\
&=\sum_{\mbox{\tiny$\begin{array}{c} y_n \geq 0\\ k=1,2,...,n\end{array}$}}\biggl\{\prod_{1\leq r<s\leq n}\biggl[\frac{1-(x_r/x_sq^{y_r-y_s})}{1-(x_r/x_s)}\biggl]\prod_{r,s=1}^n(q\frac{x_r}{x_s};q)^{-1}_{y_r}\prod_{i=1}^n(x_i)^{ny_i-(y_1+...+y_n)}(-1)^{(n-1)(y_1+...+y_n)}\\
&\quad\times q^{y_2+2y_3+...+(n-1)y_n+(n-1)[{y_1\choose 2}+...+{y_n\choose 2}]-e_2(y_1,...,y_n)}(b;q)_{y_1+...+y_n}\mathbb{T}(r,s,t,u,v,yD_z)\{z^{y_1+...+y_n}\},
\end{align*}
which is the left-hand side of \eqref{thm8_1} by \eqref{23}. The proof of Theorem \ref{thm8} is complete.
\end{proof}

\section{Generalization of Ramanujan's integrals }\label{section6}
The following integral of Ramanujan \cite{2} are quite famous.
\begin{pro}[{\cite[Eq. (2)]{2}}]\label{lem7}
For $0<q=e^{-2k^2}<1$ and $m\in\mathbb{R}$. Suppose that $\abs{abq}<1$, we have
\begin{align}\label{lem7_1}
\int^\infty_\infty\frac{e^{-x^2+2mx}}{(aq^{1/2}e^{2ikx},bq^{1/2}e^{-2ikx};q)_\infty}\td x=\sqrt{\pi}e^{m^2}\frac{(-aqe^{2mki},-bqe^{-2mki};q)_\infty}{(abq;q)_\infty}.
\end{align}
\end{pro}
In this section, we have the following generalization of Ramanujan's integrals.
\begin{thm}\label{thm9}
For $0<q=e^{-2k^2}<1$ and $m\in\mathbb{R}$. Suppose that $\abs{abq}<1$, we have
\begin{multline}\label{thm9_1}
\int^\infty_\infty\frac{e^{-x^2+2mx}}{(aq^{1/2}e^{2ikx},bq^{1/2}e^{-2ikx};q)_\infty}{}_3\phi_2\Biggl[\begin{matrix}
\begin{array}{r}
r,s,t;\\\\
u,v;
\end{array}
\end{matrix}q;yq^{1/2}e^{2ikx}\Biggr]\td x\\
=\sqrt{\pi}e^{m^2}\frac{(-aqe^{2mki},-bqe^{-2mki};q)_\infty}{(abq;q)_\infty}{}_4\phi_3\Biggl[\begin{matrix}
\begin{array}{r}
r,s,t,e^{2mki}/b;\\\\
u,v,-aqe^{2mki};
\end{array}
\end{matrix}q;ybq\Biggr].
\end{multline}
\end{thm}
\begin{rem}
For $y=0$ in Theorem \ref{thm9}, equation \eqref{thm9_1} reduces to \eqref{lem7_1}.
\end{rem}
\begin{proof}[Proof of Theorem \ref{thm9}]
Denoting the right-hand side of \eqref{thm9_1} equivalently by $f(r,s,t,u,v,a,y)$. $f(r,s,t,u,v,a,y)$ is analytic near $(r,s,t,u,v,a,y)$ and we can check that $f(r,s,t,u,v,a,y)$ satisfied \eqref{thm1_1}, so we have
\begin{align*}
f(r,s,t,u,v,a,y)=\sum_{k=0}^\infty\mu_n\phi_n^{r,s,t\choose u,v}(a,y|q),
\end{align*}\\
and
\begin{align*}
f(r,s,t,u,v,a,0)&=\sum_{k=0}^\infty\mu_na^n=\sqrt{\pi}e^{m^2}\frac{(-aqe^{2mki},-bqe^{-2mki};q)_\infty}{(abq;q)_\infty}\quad\, by\,\,\,\eqref{lem7_1}\\
&=\int^\infty_\infty\frac{e^{-x^2+2mx}}{(aq^{1/2}e^{2ikx},bq^{1/2}e^{-2ikx};q)_\infty}\td x\\
&=\int^\infty_\infty\frac{e^{-x^2+2mx}}{(bq^{1/2}e^{-2ikx};q)_\infty}\left\{\sum_{n=0}^\infty\frac{(q^{1/2}e^{2ikx})^n}{(q;q)_n}a^n\right\}\td x.
\end{align*}\par
So we have
\begin{align*}
f(r,s,t,u,v,a,y)=\int^\infty_\infty\frac{e^{-x^2+2mx}}{(bq^{1/2}e^{-2ikx};q)_\infty}\left\{\sum_{n=0}^\infty\phi_n^{r,s,t\choose u,v}(a,y|q)\frac{(q^{1/2}e^{2ikx})^n}{(q;q)_n}\right\}\td x,
\end{align*}\\
which is equal to the left-hand side of equation \eqref{thm9_1}. The proof of Theorem \ref{thm9} is complete.
\end{proof}

\section{ Concluding Remarks and Observations}
\label{conclusion}
In our present investigation, we have introduced a set of 
two $q$-operators \break $\mathbb{T}(a,b,c,d,e,yD_x)$ and 
$\mathbb{E}(a,b,c,d,e,y\theta_x)$ with to applying them
to  generalize Arjika's recently results \cite{10}, and derive transformational identities by means of the $q$-difference equations. We have   also  derived   
$U(n+1)$-type generating functions and Ramanujan's integrals involving general Al-Salam-Carlitz polynomials by means of the $q$-difference equations.

It is believed that the $q$-series and $q$-integral identities, which we
have presented in this paper, as~well as the various related recent works
cited here, will provide encouragement and motivation for further researches
on the topics that are dealt with and investigated in this paper.

In conclusion, we find it to be worthwhile to remark that some potential
further applications of the methodology and findings, which we have
presented here by means of the $q$-analysis and the $q$-calculus, 
can be found in the study of
the zeta and $q$-zeta functions as well as their related 
functions of Analytic Number Theory
(see, for example, \cite{HMS-JAEC2019, HMS-PIMM2019}; see also
\cite{Srivastava2012}) and also in the study of analytic and
univalent functions of Geometric Function Theory {\it via} 
number-theoretic entities (see, for example, \cite{HMS-Symmstry2020}).

\medskip

\noindent
{\bf Conflicts of Interest:} The authors declare that they 
have no conflicts of interest.\\
{\bf Availability of data and material:} Not applicable. \\
 {\bf Funding:} This work was supported by the Zhejiang Provincial 
Natural Science Foundation of China (No. LY21A010019). \\
 {\bf Authors' contributions:} Both authors equally contributed to this manuscript   and approved the final version.  \\
 {\bf Competing interests:}  The authors declare that there are no competing interests.  \\
 {\bf Acknowledgements:} Not applicable.

\end{document}